OpenStudy (anonymous):

Equations that are not equal 1/25

OpenStudy (thomas5267):

1+1?

OpenStudy (freckles):

what does that mean

OpenStudy (neonumbrella5115):

Are there any choices??

OpenStudy (solomonzelman):

\(\large\color{blue}{ \displaystyle \frac{ 1 }{25} }\) \(\large\color{green}{ \displaystyle\left( \frac{ 1 }{5}\right) ^2 }\) \(\large\color{blue}{ \displaystyle 5^{-2} }\) \(\large\color{green}{ \displaystyle 0.04 }\) \(\large\color{blue}{ \displaystyle 4\text{%} }\)